A water-tight open box 11.19 feet in length, 5.28 feet wide, and 6.16 feet high weighs 8330 pounds.

If this "ship" is placed in water, how much of it will remain above water, in feet?
Answer #1

ANO - m=8330 pounds = 3778.4 kg A - L Xho SF=F6-MG=0 FG = ng = 3718.4 (5.28 X0.3048m) (11.1980.3048 m) (1000h) - 3778.4 h = 0